2. Calculating tax llabilities Calculating Taxable Income Taxable income $1 $9,075 $9,076-$36,900 $36,901-$89,350 $89,351-$186,350 $186,351-$405,100 $405,101-$406,750 Over $406,751 Rate 10.0% 15.0% 25.0% 28.0% 33.0% 35.0% 39.6% Using the tax table shown at the left, compute the tax liability for each individual in the scenarios presented, rounding the liability to the nearest dollar. In addition, use the dropdown lists to identify the marginal tax rate and average tax rate for both Tracy and Ben. Tracys Tax Scenario Tracy is a young professional who lives in downtown Chicago. She has taxable income of $94,735 as an advertising account executive for Taggart Goods Corp What is Tracys tax liability? What is Tracys top marginal tax rate? What is Tracys average tax rate? Bens Tax Scenario Ben is a young professional who lives in downtown Atlanta. He has taxable income of $63,500 as a medical sales rep for Tull Cybermatic Corp. What is Bens tax liability? s What is Bens top marginal tax rate? what is Bens average tax rate? Grade It Now
Answer #1

Tracy:

Taxable Income = $ 94735

Total Tax Liability is calculated by adding the tax expense for each slab that the taxable income crosses, followed by adding up these individual expenses to arrive at the final result. The same is demonstrated below:

Taxable Income: $ 1 - $ 9075 and Tax Rate = 10 %

Tax Liability = 0.1 x 9075 = $ 907.5

Taxable Income: $ 9076 - $ 36900 and Tax Rate = 15 %

Tax Liability = (36900 - 9075) x 0.15 = $ 4173.75

Taxable Income: $ 36901 - $ 89350 and Tax Rate = 25 %

Tax Liability = (89350 - 36900) x 0.25 = $ 13112.5

Taxable Income: $ 89351 - $ 186350 and Tax Rate = 28 %

Tax Liability = (94735 - 89350) x 0.28 = $ 1507.8 (this interval is not used up entirely as Tracy's taxable income does not cross the upper limit of this interval)

Total Tax Liability = 907.5 + 4173.75 + 13112.5 + 1507.8 = $ 19701.55

Marginal Tax Rate is the Tax Rate corresponding to the Tax Bracket (Interval) in which Tracy's taxable income lies which in this case is the $ 89351 to $ 186350 bracket. Hence, the marginal tax rate is 28%

Average Tax Rate = (Total Tax Liability / Taxable Income) = (19701.55 / 94735) ~ 0.20796 or 20.796 %

Ben:

Taxable Income = $ 63500

Total Tax Liability is calculated by adding the tax expense for each slab that the taxable income crosses, followed by adding up these individual expenses to arrive at the final result. The same is demonstrated below:

Taxable Income: $ 1 - $ 9075 and Tax Rate = 10 %

Tax Liability = 0.1 x 9075 = $ 907.5

Taxable Income: $ 9076 - $ 36900 and Tax Rate = 15 %

Tax Liability = (36900 - 9075) x 0.15 = $ 4173.75

Taxable Income: $ 36901 - $ 89350 and Tax Rate = 25 %

Tax Liability = (63500 - 36900) x 0.25 = $ 6650 (this interval is not used up entirely as Ben's taxable income does not cross the upper limit of this interval)

Total Tax Liability = 907.5 + 4173.75 + 6650 = $ 11731.25

Marginal Tax Rate is the Tax Rate corresponding to the Tax Bracket (Interval) in which Tracy's taxable income lies which in this case is the $ 369001 to $ 89350 bracket. Hence, the marginal tax rate is 25 %

Average Tax Rate = (Total Tax Liability / Taxable Income) = (11731.25 / 63500) ~ 0.18474 ~ 18.474 %